'wo identical thin rings, each ofradiuses X are coaxially placed at a distance,ll apart. lt e1 and p2 are respectively the charges uniformly spread on the two rings. The work done in moving a charge g fronr the centre ofone ring to that ofthe second ring is

Grade:12th pass

1 Answers

Arun
4 years ago

E and V at a point P that lies on the axis of ring -

\dpi{100} E_{x}=\frac{kQx}{\left ( x^{2}+R^{2} \right )^{\frac{3}{2}}}  ,   V=\frac{kQ}{\left ( x^{2}+R^{2} \right )^{\frac{1}{2}}}

-

 

 

Charged Circular ring -

A charged circular ring of radius R and charge Q.

- wherein

 

 

 

 

Potential at the centre of first ring V_{A}=\frac{Q_{1}}{4\pi \varepsilon _{0}R}+\frac{Q_{2}}{4\pi \varepsilon _{0}\sqrt{R^{2}+R^{2}}}

Potential at the centre of second ring V_{B}=\frac{Q_{2}}{4\pi \varepsilon _{0}R}+\frac{Q_{1}}{4\pi \varepsilon _{0}\sqrt{R^{2}+R^{2}}}

Potential difference between the two centres V_{A}-V_{B}=\frac{\left ( \sqrt{2}-1 \right )\left ( Q_{1}-Q_{2} \right )}{4\pi \varepsilon _{0}R\sqrt{2}}

\therefore Work\; done     W=\frac{q\left ( \sqrt{2}-1 \right )\left ( Q_{1}-Q_{2} \right )}{4\pi \varepsilon _{0}R\sqrt{2}}